OpenStudy (anonymous):

How do I find the equation for a line on a graph? Either an explanation or a link to an explanation would be appreciated! Thanks.

jimthompson5910 (jim_thompson5910):

Do you have the graph?

OpenStudy (anonymous):

No, but I have two points. (8,3) and (4,3)

jimthompson5910 (jim_thompson5910):

what is the slope of the line through those points?

OpenStudy (anonymous):

Erm... heh. I don't know? Sorry...

jimthompson5910 (jim_thompson5910):

you would use the slope formula \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

jimthompson5910 (jim_thompson5910):

(x1,y1) is your first point (x2,y2) is your second point

OpenStudy (anonymous):

So m=0/-4?

jimthompson5910 (jim_thompson5910):

which reduces to 0

jimthompson5910 (jim_thompson5910):

the slope of m = 0 means that we have a completely flat horizontal line

jimthompson5910 (jim_thompson5910):

now we turn to y = mx+b

jimthompson5910 (jim_thompson5910):

m = 0

jimthompson5910 (jim_thompson5910):

x and y are from either point you pick so let's say you pick (8,3) that means x = 8 and y = 3

jimthompson5910 (jim_thompson5910):

so... y = mx+b 3 = 0*8 + b b = ???

OpenStudy (anonymous):

b=3

OpenStudy (anonymous):

Right?

jimthompson5910 (jim_thompson5910):

m = 0 and b = 3 y = mx+b y = 0x + 3 y = 3

jimthompson5910 (jim_thompson5910):

so the final answer is y = 3

OpenStudy (anonymous):

Okay, thank you! :)

jimthompson5910 (jim_thompson5910):

you're welcome